OpenStudy (anonymous):

really need HELP!!! The formula for that area of a triangle is A = 1/2}bh. If the base is 5\sqrt 3 and the height is \sqrt {18} , what is the area of this triangle in radical form?

OpenStudy (campbell_st):

\[Area = \frac{1}{2} \times \frac{5}{\sqrt{3}} \times \sqrt{18} = \frac{1}{2} \times 5 \times \sqrt{\frac{18}{3}}\] simplify the fraction inside the radical for the answer.

OpenStudy (mathstudent55):

@campbell_st Perhaps the poster meant this: \(A = \dfrac{bh}{2} \) \(A = \dfrac{5 \sqrt{3} \times \sqrt{18} }{2} \)

OpenStudy (campbell_st):

well if thats the case its \[Area = \frac{5 \times \sqrt{3 \times 18}}{2}\]

OpenStudy (anonymous):

I dont have a calculator to find the sqaure root can someone do it please ?

OpenStudy (swissgirl):

\(Area = \large \frac{5* \sqrt{54}}{2} = \frac{5* \sqrt{9*6}}{2}=\frac{15\sqrt{6}}{2} \) If you would not like the sqrt in your answer thennn \( \sqrt{6}\) is approximately 2.45 sooo \( Area = \large \frac{15*2.45}{2}\)

OpenStudy (anonymous):

so your answer would be 18.375 ?

OpenStudy (swissgirl):

yess aproximately

OpenStudy (anonymous):

is this in radical form ??

OpenStudy (swissgirl):

nooooooo Radical form is what I showed above \( Area = \large \frac{15\sqrt{6}}{2}\)

OpenStudy (anonymous):

what is the area of this triangle in radical form? in the question it says this so how would i write this ?

OpenStudy (anonymous):

for the answer

OpenStudy (swissgirl):

There I showed u above

OpenStudy (anonymous):

ok so can u tell me the final answer ?

OpenStudy (anonymous):

i dont have that sqrt sign on my keyboard